# KeyDiff: Key Similarity-Based KV Cache Eviction for Long-Context LLM Inference in Resource-Constrained Environments

### v3

@@ Appendix E Math-500 reasoning benchmark
Table 3 summarizes the Math-500 evaluation results for Llama-8B. As shown in the table, token eviction methods generally perform well even with KV cache budgets that are strictly smaller than the maximum generation length. Surprisingly, we found that KeyDiff slightly outperforms baseline methods in certain configurations (e.g., KeyDiff with a 2K budget for 4K generation length, and KeyDiff with a 4K budget for 8K generation length).
To further analyze the effectiveness of token eviction, we measure accuracy in cases where the context length of the baseline method (i.e., prompt length + generation length) exceeds the available KV cache budget. As shown in Table 4, for samples where eviction is actively triggered, KeyDiff continues to outperform the token eviction baseline (SnapKV), and often achieves accuracy close to or slightly better than the non-evicting baseline.
We also conducted a similar evaluation on DeepSeek-R1-Distill-Qwen-7B and observed a slight performance degradation for token eviction methods compared to full KV cache baselines (See Tables 5 and 6.) However, KeyDiff still demonstrates comparable performance to SnapKV overall. This discrepancy may stem from architectural differences that Llama uses a lower GQA [2] ratio than Qwen, which results in more information compression in the KV cache. We hypothesize that models with more compression like Qwen are more sensitive to eviction since each evicted token contains more information in Qwen than Llama by design.
@@
Max Gen. Length $=$ 4K
Num Tokens $>$ 1K (497/500 samples)
@@
Max Gen. Length $=$ 8K
Num Tokens $>$ 2K (353/500 samples)
@@
Max Gen. Length $=$ 32K
Num Tokens $>$ 2K (353/500 samples)
@@
Max Gen. Length $=$ 4K
Num Tokens $>$ 1K (497/500 samples)
@@
Max Gen. Length $=$ 8K
Num Tokens $>$ 2K (317/500 samples)
@@
Max Gen. Length $=$ 32K
Num Tokens $>$ 2K (360/500 samples)
@@ F.1 Empirical Motivation for KeyDiff Setup
To generate Figure 11, we used the first sample from the test split of the narrativeqa task in LongBench to prefill the KV cache of Llama3.2-3B-Instruct with a block size of $B=128$. The KV cache had a maximum size of $4096$ while the sample was much longer, requiring KV eviction. We applied PCA to the key cache and repeated the process for sink attention, TOVA and KeyDiff.
To construct Figure 12, we sample 100 prompts from the Qasper dataset in LongBench [4], compute the log determinant of $KK^{T}$ of the keys in the KV caches of each head and layer of Llama 3.2-3B-Instruct using a cache budget of $N=2048$ and a block size of $B=128$, and plot the distribution in Figure 12. We show this key distribution for sink attention, TOVA and KeyDiff.
@@ F.3 Longbench dataset statistics
We measure the number of tokens in the samples using LLama tokenizer [31]. As shown in Figure 18, LongBench exhibits variability in sample length from the datasets.
@@
Average Compression $\text{Average Compression Ratio}=\frac{1}{I}\sum_{i=1}^{I}\frac{N}{L_{i}},$
@@
As summarized in Table 8, 2K cache budgets have a 0.31 average compression ratio, which indicates 69% of input prompts are compressed. Our largest cache budget, 8K, exhibits a 0.77 average compression ratio.
@@ F.4 Additional Results
Table 9: Resource unrestricted LongBench Results (Higher is better). All methods processes the input prompt in parallel (i.e., block size =absent=$\infty$= ∞) and make an eviction decision with all token information in the input. The token eviction is made at every step of generation if the budget exceeds. We highlight the methods showing the best performance within a given budget with boldface. We omit NarrativeQA from evaluation due to higher chance of OOM errors.
@@
[…]
@@ Appendix G Ablation study
Table 14: Anchor vector ablation study. Average of Full Longbench results for Llama 3.2-3B-Instruct. KeyDiff results match Table 10. (Higher is better)
@@
We have mainly evaluated KeyDiff using the method described in Equation 7. Scores to determine eviction are measured via cosine similarity with an anchor vector which can be computed in several ways. We run LongBench on Llama3.2-3B-Instruct with eviction policies using the following anchor choices: pairwise cosine similarity from Equation 7, denoted $\operatorname{Pairwise}$; KeyDiff, using the mean of all normalized keys as an anchor, and using the median of keys as an anchor, denoted $\operatorname{Median}$. Table 14 summarizes the average LongBench accuracy for the different methods to Llama 3.2-3B-Instruct. KeyDiff shows similar average scores to $\operatorname{Pairwise}$. Additionally, KeyDiff and $\operatorname{Median}$ show similar scores, demonstrating that KeyDiff is robust to the selection of the anchor design.
Table 15: Distance metric ablation study. Average of Full Longbench results of Llama 3.2-3B-Instruct. KeyDiff results match Table 10. (Higher is better)
@@
We use cosine similarity as the scoring metric for eviction in KeyDiff based on our discussion in Section 2.2. This could be replaced with other metrics like the dot product or Euclidean distance. We evaluate KeyDiff variants using dot product and Euclidean distance as the similarity metric, denoted $\operatorname{DotProd}$ and $\operatorname{Euclidean}$ respectively, and report the results in Table 15. KeyDiff and $\operatorname{DotProd}$ show similar performance for 6K and 8K budgets. However, KeyDiff outperforms $\operatorname{DotProd}$ for smaller cache sizes. This implies that considering both the direction and the magnitude of the keys to compute similarity are important for identifying the tokens to evict. On the other hand, $\operatorname{Euclidean}$ shows a significant performance drop relative to KeyDiff.

### v4

@@ Appendix E Math-500 reasoning benchmark
Table˜4 summarizes the Math-500 evaluation results for Llama-8B. As shown in the table, token eviction methods generally perform well even with KV cache budgets that are strictly smaller than the maximum generation length. Surprisingly, we found that KeyDiff slightly outperforms baseline methods in certain configurations (e.g., KeyDiff with a 2K budget for 4K generation length, and KeyDiff with a 4K budget for 8K generation length).
To further analyze the effectiveness of token eviction, we measure accuracy in cases where the context length of the baseline method (i.e., prompt length + generation length) exceeds the available KV cache budget. As shown in Table˜5, for samples where eviction is actively triggered, KeyDiff continues to outperform the token eviction baseline (SnapKV), and often achieves accuracy close to or slightly better than the non-evicting baseline.
We also conducted a similar evaluation on DeepSeek-R1-Distill-Qwen-7B and observed a slight performance degradation for token eviction methods compared to full KV cache baselines (See Tables˜6 and 7.) However, KeyDiff still demonstrates comparable performance to SnapKV overall. This discrepancy may stem from architectural differences that Llama uses a lower GQA [2] ratio than Qwen, which results in more information compression in the KV cache. We hypothesize that models with more compression like Qwen are more sensitive to eviction since each evicted token contains more information in Qwen than Llama by design.
@@
Max Gen. Length == 4K
Num Tokens >> 1K (497/500 samples)
@@
Max Gen. Length == 8K
Num Tokens >> 2K (353/500 samples)
@@
Max Gen. Length == 32K
Num Tokens >> 2K (353/500 samples)
@@
Max Gen. Length == 4K
Num Tokens >> 1K (497/500 samples)
@@
Max Gen. Length == 8K
Num Tokens >> 2K (317/500 samples)
@@
Max Gen. Length == 32K
Num Tokens >> 2K (360/500 samples)
@@ F.1 Empirical Motivation for KeyDiff Setup
To generate Figure˜11, we used the first sample from the test split of the narrativeqa task in LongBench to prefill the KV cache of Llama3.2-3B-Instruct with a block size of $B=128$. The KV cache had a maximum size of $4096$ while the sample was much longer, requiring KV eviction. We applied PCA to the key cache and repeated the process for sink attention, TOVA and KeyDiff.
To construct Figure˜12, we sample 100 prompts from the Qasper dataset in LongBench [4], compute the log determinant of $KK^{T}$ of the keys in the KV caches of each head and layer of Llama 3.2-3B-Instruct using a cache budget of $N=2048$ and a block size of $B=128$, and plot the distribution in Figure˜12. We show this key distribution for sink attention, TOVA and KeyDiff.
@@ F.3 Longbench dataset statistics
We measure the number of tokens in the samples using LLama tokenizer [31]. As shown in Figure˜18, LongBench exhibits variability in sample length from the datasets.
@@
$\text{Average Compression Ratio}=\frac{1}{I}\sum_{i=1}^{I}\frac{N}{L_{i}},$
@@
As summarized in Table˜9, 2K cache budgets have a 0.31 average compression ratio, which indicates 69% of input prompts are compressed. Our largest cache budget, 8K, exhibits a 0.77 average compression ratio.
@@ F.4 Additional Results
Table 10: Resource unrestricted LongBench Results (Higher is better). All methods processes the input prompt in parallel (i.e., block size $\infty$) and make an eviction decision with all token information in the input. The token eviction is made at every step of generation if the budget exceeds. We highlight the methods showing the best performance within a given budget with boldface. We omit NarrativeQA from evaluation due to higher chance of OOM errors.
@@
Table 14: KeyDiff + Recent tokens Llama-3.1-8B/3.2-3B-Instruct LongBench Results with $B=128$ (Higher is better). We highlight the methods showing the best performance within a given budget with boldface. $\text{X}\%$ indicates $\text{X}\%$ of KV cache budget is reserved to keep the recent tokens, while the remaining cache budget is managed by KeyDiff algorithm. †: A subset of samples were evaluated due to OOM errors (183/200 samples are evaluated).
@@ Appendix G Ablation study
Table 15: Anchor vector ablation study. Average of Full Longbench results for Llama 3.2-3B-Instruct. KeyDiff results match Table˜11. (Higher is better)
@@
We have mainly evaluated KeyDiff using the method described in Equation˜7. Scores to determine eviction are measured via cosine similarity with an anchor vector which can be computed in several ways. We run LongBench on Llama3.2-3B-Instruct with eviction policies using the following anchor choices: pairwise cosine similarity from Equation˜7, denoted $\operatorname{Pairwise}$; KeyDiff, using the mean of all normalized keys as an anchor, and using the median of keys as an anchor, denoted $\operatorname{Median}$. Table˜15 summarizes the average LongBench accuracy for the different methods to Llama 3.2-3B-Instruct. KeyDiff shows similar average scores to $\operatorname{Pairwise}$. Additionally, KeyDiff and $\operatorname{Median}$ show similar scores, demonstrating that KeyDiff is robust to the selection of the anchor design.
Table 16: Distance metric ablation study. Average of Full Longbench results of Llama 3.2-3B-Instruct. KeyDiff results match Table˜11. (Higher is better)
@@
We use cosine similarity as the scoring metric for eviction in KeyDiff based on our discussion in Section˜2.2. This could be replaced with other metrics like the dot product or Euclidean distance. We evaluate KeyDiff variants using dot product and Euclidean distance as the similarity metric, denoted $\operatorname{DotProd}$ and $\operatorname{Euclidean}$ respectively, and report the results in Table˜16. KeyDiff and $\operatorname{DotProd}$ show similar performance for 6K and 8K budgets. However, KeyDiff outperforms $\operatorname{DotProd}$ for smaller cache sizes. This implies that considering both the direction and the magnitude of the keys to compute similarity are important for identifying the tokens to evict. On the other hand, $\operatorname{Euclidean}$ shows a significant performance drop relative to KeyDiff.
@@ G.1 Needle in a Haystack results for Sink attention
Appendix H Additional Experiments and Analyses
To further substantiate the empirical findings presented in Sections˜3 and 4, we provide an extended set of experiments and complementary analyses. These additional studies examine the correlation between key geometry and attention, evaluate retrieval-critical and reasoning scenarios, and measure the efficiency of KeyDiff under constrained hardware settings. Together, they offer a broader view of the method’s robustness, efficiency, and general applicability across diverse inference conditions.
H.1 Retrieval-Critical Evaluation: Phonebook Lookup
We evaluate KeyDiff on a retrieval-critical setting using the Phonebook Lookup task, where the model retrieves a phone number corresponding to a queried name from a long list of entries. Accuracy is averaged across five random phonebooks of varying lengths. As shown in Table˜17, KeyDiff maintains high retrieval accuracy for shorter contexts and degrades more gracefully than attention-based baselines as context length increases.
Table 17: Accuracy (%) on the Phonebook Lookup task using Llama-3.2-3B-Instruct with a 6k cache budget.
Method 100 478 856 1233 1611 1989 2367 2744 3122 3500
Dense 1.0 1.0 1.0 0.8 1.0 0.8 0.6 0.6 0.8 0.6
KeyDiff 1.0 1.0 1.0 0.6 0.4 0.2 0.0 0.0 0.2 0.2
TOVA 1.0 1.0 1.0 0.4 0.2 0.0 0.0 0.0 0.0 0.0
H.2 RULER Benchmark Validation
To assess the generalizability of KeyDiff across diverse architectures, we reference the results reported by the community-maintained KVPress RULER benchmark. KeyDiff consistently achieves competitive or superior leaderboard scores compared to methods such as TOVA, SnapKV, QFilter, and Knorm on both Llama-3.2-3B and Qwen-3-8B backbones, demonstrating its robustness and transferability.
H.3 Needle-in-a-Haystack Recall Saturation
We further analyze the recall behavior of KeyDiff under varying context lengths in the Needle-in-a-Haystack benchmark. Table 18 reports the recall difference between KeyDiff and the full-cache baseline as a function of depth and context length. As shown in Table˜18, KeyDiff achieves near-parity recall with the full-cache baseline up to 30k tokens, confirming its stability under long-context compression.
Table 18: Recall difference (KeyDiff-Full) on NIAH benchmark for Llama-3.2-3B.
Depth 1000 4222 7444 10667 13889 17111 20333 23556 26778 30000
0.0 0.02 -0.06 -0.06 0.20 0.16 -0.16 0.18 0.20 0.06 0.04
11.0 0.10 -0.04 -0.04 0.30 0.28 -0.18 0.26 0.28 0.10 -0.08
22.0 0.04 -0.12 0.00 0.32 0.30 -0.18 0.16 0.28 0.14 -0.04
33.0 0.02 0.06 -0.20 0.34 0.08 -0.16 0.26 0.26 0.20 -0.16
44.0 -0.06 -0.08 -0.16 0.30 0.26 -0.12 0.24 0.22 -0.10 0.04
56.0 -0.06 0.00 -0.24 0.26 0.26 -0.20 0.18 0.14 0.04 -0.02
67.0 0.02 0.02 -0.14 0.36 0.28 -0.26 0.24 0.16 0.04 -0.08
78.0 -0.10 0.08 -0.26 0.26 0.30 -0.26 0.20 0.30 0.02 -0.04
89.0 0.02 0.08 -0.06 0.30 0.26 -0.32 0.30 0.22 0.06 -0.12
100.0 0.00 -0.06 -0.26 0.00 -0.06 -0.02 0.00 -0.04 0.14 -0.04
H.4 On-Device Latency Evaluation
We further evaluate the runtime characteristics of KeyDiff on a mobile-class device by measuring the latency required to compute key-eviction scores on an Android platform. All methods are tested on a recent Samsung smartphone under identical FP16 inference precision, and the results are normalized by the latency of KeyDiff with a cache budget of 512 KVs.
As shown in Table˜19, KeyDiff performs on par with competing methods for small cache sizes and achieves substantially lower scoring latency as the cache size increases, demonstrating both scalability and minimal overhead on edge hardware. Minor runtime fluctuations can be attributed to kernel-level optimizations and proprietary hardware characteristics.
Table 19: Relative latency of key scoring on an Android device (normalized by KeyDiff at 512 KVs). Lower is better.
Method 512 1024 2048 4096 8192
SnapKV 1.92 2.46 4.50 8.94 11.42
TOVA 1.02 0.94 1.25 1.95 3.06
KeyDiff 1.00 1.03 1.01 1.07 1.37
H2O 1.10 0.99 1.52 2.32 4.23
NeurIPS Paper Checklist
  1. 1.
    Claims
    • Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
    • Answer: [Yes]
    • Justification: The claims made in the abstract are corroborated by both our theoretical justification Section˜3.3, experiments Section˜4 and auxiliary results and discussion.
  2. 2.
    Limitations
    • Question: Does the paper discuss the limitations of the work performed by the authors?
    • Answer: [Yes]
    • Justification: In Section˜6 we discuss the fact that KeyDiff is only tested on decoder based models featuring GQA and would like to extend it in the future to models with different designs such as MLA.
  3. 3.
    Theory Assumptions and Proofs
    • Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
    • Answer: [Yes]
    • Justification: Proofs of all results are discussed briefly in their statement in Section˜3.3 and rigorously proved in Section˜C.3.
  4. 4.
    Experimental Result Reproducibility
    • Question: If the contribution is a dataset or model, what steps did you take to make your results reproducible or verifiable? Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), release of a model checkpoint, or other means that are appropriate to your research.
    • Answer: [N/A]
    • Justification: The paper releases neither a dataset or a model, though the cache management algorithm can be easily implemented using the description from the paper.
  5. 5.
    Open access to data and code
    • Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
    • Answer: [No]
    • Justification: Code is not provided as it is proprietary. However the algorithm may be readily implemented and results replicated using the description in the paper.
  6. 6.
    Experimental Setting/Details
    • Question: If you ran experiments, did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? The full details can be provided with the code, but the important details should be in the main paper, and information about how hyperparameters were selected should appear either in the paper or supplementary materials.
    • Answer: [Yes]
    • Justification: All hyperparameters selected for the benchmarks run to evaluate KeyDiff are specified in Sections˜4, F.2 and E. Hyperparameters to recreate figures are given in figure descriptions.
  7. 7.
    Experiment Statistical Significance
    • Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
    • Answer: [Yes]
    • Justification: The only experiment using stochasticity Figure˜7 and Figure˜17 reports error bars and statistical significance. The Math500 results as well discuss the use of sampling and stochasticity in generating responses. Aside from this, the remaining results are deterministic.
  8. 8.
    Experiments Compute Resources
    • Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
    • Answer: [Yes]
    • Justification: In all relevant cases, particularly experiments on TTFT Figure˜17 and Figure˜7, we mention the GPUs and compute resources used.
  9. 9.
    Code Of Ethics
    • Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
    • Answer: [Yes]
    • Justification: The research proposes a new cache management algorithm based on the geometry of the attention mechanism in decoder-based LLMs. It does not involve human subjects or obviously ethically sensitive applications, and it conforms to the NeurIPS Code of Ethics.
  10. 10.
    Broader Impacts
    • Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
    • Answer: [N/A]
    • Justification: This research is foundational in nature and is not tied to particular applications or deployments. As such we do not see any direct path to negative applications of our work.
  11. 11.
    Safeguards
    • Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
    • Answer: [N/A]
    • Justification: The paper proposes a new cache eviction algorithm. Neither the models used, data or cache eviction algorithm appear to pose a high risk for abuse necessitating release safeguards beyond standard practices.
  12. 12.
    Licenses for existing assets
    • Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
    • Answer: [Yes]
    • Justification: Creators for datasets/benchmarks LongBench and Math500 are properly credited with citations. Open source models from the Llama and Qwen model classes are used to test KeyDiff are properly cited as are their deepseek distill r1 variants.
  13. 13.
    Assets
    • Question: If you are releasing new assets, did you document them and provide these details alongside the assets? Researchers should communicate the details of the dataset or the model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • Answer: [N/A]
    • Justification: The research does not release new assets.
  14. 14.
    Crowdsourcing and Research with Human Subjects
    • Question: If you used crowdsourcing or conducted research with human subjects, did you include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)? Including this information in the supplemental material is fine, but if the main contribution of your paper involves human subjects, then we strongly encourage you to include as much detail as possible in the main paper. According to the NeurIPS Code of Ethics, you must pay workers involved in data collection, curation, or other labor at least the minimum wage in your country.
    • Answer: [N/A]
    • Justification: The research does not involve crowdsourcing experiments or research with human subjects.
  15. 15.
    IRB Approvals
    • Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
    • Answer: [N/A]
    • Justification: The research does not involve human subjects, therefore IRB approval is not applicable.
  16. 16.
    Declaration of LLM usage
    • Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
    • Answer: [N/A]
    • Justification: The research does not include any input from an LLM. In particular the core methodology and research is original to the authors.